The value of q for which the difference between the roots of the equation x^2-qx+8=0 is 2 are
+-2
+-4
+-6
+-8

Found 2 solutions by ikleyn, greenestamps:
Answer by ikleyn(52778)

.
I easily can guess:  the roots 2 and 4 gives the product of 8 (the constant term) and the difference of 2.


Their sum is 6, which should be " q", according to Vieta's theorem.


So, the answer is q = 6, based on my guessing.



Let's look what the Algebra solution will give us.


Let x and (x-2) are the roots.

Then their product is 8:

    x*(x-2) = 8

    x^2 - 2x - 8 = 0

    (x-4)*(x+2) = 0.


So, there are 2 roots:  x= 4  and  x= -2.


The value  x= 4  gives that two roots  4 and 2 which I guessed above, with the value  of q= 6.


The value x= -2 gives two roots  -2 and -4, with the value of q = -6.


So, the problem has two answers:  q= 6  and  q= -6.    (Third line of the answers' choice)

Thus Algebra solution helped me to find 2 answers to the problem question: more than I could guess (!)


Answer by greenestamps(13198)



According to Vieta's Theorem, for the given equation x^2-qx+8=0, the product of the roots is 8 and the sum of the roots is q.

Then, given that the difference between the two roots is 2, a very little bit of mental arithmetic (product of two numbers = 8; difference = 2) finds two solutions -- roots of 2 and 4, or roots of -2 and -4.

And those roots make q either 6 or -6.

ANSWER: The third choice, -6, is ONE OF the values for q for which the difference between the two roots of the equation is 2.